OpenStudy (bananas1234):

divide 12mn - 30m ------------ 6m

OpenStudy (skullpatrol):

any ideas? :-)

OpenStudy (bananas1234):

not sure, sorry :/

OpenStudy (skullpatrol):

factor out the "m" from the numerator first

OpenStudy (bananas1234):

how do i factor it out?

OpenStudy (skullpatrol):

recall: ab + ac = a( b + c )

OpenStudy (skullpatrol):

a is said to be factored out^

OpenStudy (bananas1234):

soo 6m + m6 = m(6) i think that is probably wrong...

OpenStudy (skullpatrol):

try again using $$\Huge ab + ac = a( b +c )$$ as a model :-)

OpenStudy (skullpatrol):